Find (i') the lateral or curved surface area of a closed cylind rical petrol storage tank that is 4.2 m in diameter and 4.5 m high. (ii) how much steel was actually used, if of the steel actually used was wa.vted in ma.king the tank.

Find
(i') the lateral or curved surface area of a closed cylind rical petrol storage tank that is 4.2 m in diameter and 4.5 m high.

(ii) how much steel was actually used, if of the steel
actually used was wa.vted in ma.king the tank.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i') Lateral surface area = 2 >< 7 >< 2.1 >< 4.6 m2 = 69.4 m2• (ii) 'lbtal surface area of the tank = 69.4 + 2 >< Let total st.eel used = x m2 )( 2.1 )( 2.1 = 69.4 + 27.72 = 87.12 m2• :. (x- x) = 87.12 x = >< 87.12 m2 = 95.04 m2•